# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

### v3

@@ 7 Impossibility results: The Ten Failed Methods
    Single-Bit Flags: A boolean flag alone is non-monotone; specific steps (e.g., ‘merge void‘) can toggle the flag from 0 to 1.
@@
    Structural Depth: Simple depth counters ($\kappa$) fail on collapsing rules like merge_cancel where the result may have equal depth to the input.
@@
    Local Helper Lemmas: Attempts to patch specific rules with local lemmas failed due to the global interconnectedness of the recursion.
@@
    Unchecked Recursion: The raw rec_succ rule itself serves as the ultimate counterexample; without the SafeStep guards, it permits unbounded recursion that no internal measure can bound.
These ten failures collectively demonstrate that the termination barrier is structural: the system’s internal logic cannot capture the necessary invariants without the sophisticated machinery of the triple-lexicographic measure.
@@
More generally: any polynomial proof for a system with duplicating rules must hardcode specific constants (e.g., $M(\mathrm{void})=2$) to satisfy inequalities like $2M(s)>M(s)+1$. Such proofs import external arithmetic properties and impose arbitrary values on operators, violating the principle that operators’ semantics should be defined solely by rewrite rules. Changing the hardcoded constant (e.g., $M(\mathrm{void})=1$) collapses the proof. This structural failure is why polynomial interpretations cannot orient generic duplicating rules without external axioms.
Internally definable measures
A simple contract for internally defined termination measures structures these negative results:
Definition 1 (Internally definable measure).
An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with explicit witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t. $_{\beta}$.
@@ 8 Decidability of Reachability
Theorem 3 (Fixed-target reachability).
In a strongly normalizing TRS, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
@@
The model assumes a finite first-order TRS over finite terms. The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
@@
Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1]. This situates our “normalize and compare” decision procedure for KO7 within the established landscape.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [3], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [6].
9 Conjecture: Full Termination of Any Relational Operator-Only TRS
What “relational” means
A relational operator-only TRS is one capable of sequential computation: any system that can express “do $X$, then do YY” without external conditions. The terms are equivalent: “relational” = “capable of sequential computation.”
9.1 The Relational Nature of KO7 and Recursors
The classification of KO7 as “relational” is not arbitrary; it describes the system’s capability to internalize sequential control flow. In flat rewrite systems (e.g., group axioms), rules rewrite terms to equivalent terms. In a relational TRS like KO7, the rewrite rules define a computational process.
@@ 9.1 The Relational Nature of KO7 and Recursors
is not merely a reduction; it is a program. It applies the “step” term $s$ repeatedly, $n$ times. This is functionally identical to the transition function $\delta$ of a Turing machine or the fixed-point combinator $Y$ in lambda calculus. Both mechanisms allow a system to execute a sequence of operations of unbounded length.
@@
Any system capable of this sequential computation must have a mechanism to carry the “step” logic ($s$) forward into the next iteration. In KO7, this is achieved by duplicating $s$: it appears once in the current application ($\mathrm{app}\,s\dots$) and once in the recursive call ($\mathrm{rec}\Delta\dots s\dots$).
This duplication is the structural barrier. As demonstrated in the Impossibility Results (Section 7), this redistribution of $s$ defeats all additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). A system that internalizes its own control flow in this way cannot prove its own termination using simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
@@
Standard systems escape this trap by stepping outside the system.
  • Proof Assistants (Coq, Lean): Use a meta-logical layer (guard predicates, structural recursion checks) to enforce termination externally. The system doesn’t prove itself; the compiler proves the system.
@@
KO7’s SafeStep fragment follows this pattern by imposing external guards ($\delta$-flag). The Conjecture posits that for the full, unguarded system, no such internal proof exists.
@@
Any operator system capable of ordered computation requires a recursor. A recursor redistributes its step argument across recursive calls. This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
Definition 2 (Internally definable measure).
@@
An internally definable termination method for signature $\Sigma$ uses only: simplification orders with fixed precedence/status on $\Sigma$ (LPO/RPO/MPO), DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms. No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
Conjecture 1 (Full Termination Conjecture for Relational TRSs).
@@
No relational operator-only TRS can have its full-system termination proved by internally definable methods. Specifically: let $R$ be an operator-only TRS with a recursor rule of the form
@@
No internally definable measure (Definition 2) proves termination of $R$ when the step argument $s$ is unrestricted. The structural redistribution of $s$ into both the function application and the recursive call defeats every additive, polynomial, or path-order measure that does not import external arithmetic or ordinal axioms. Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside scope and does not count as an internal termination argument.
@@
The conjecture applies to any operator-only TRS capable of sequential computation: any system that can express “do $X$, then do YY” using only rewrite rules, without external axioms, imported arithmetic, or borrowed logic. Systems without such capability (purely ground, non-recursive, or lacking step-iteration) fall outside the conjecture’s scope. The claim is that once a system can compute sequentially without external conditions, its full termination escapes internal proof methods.
@@
In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, every additive measure fails to decrease strictly. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$. The guarded SafeStep fragment terminates because the $\delta$-phase bit restricts allowable step arguments; without such restriction, the full system admits unbounded recursive depth. The gap between internally definable ranking functions and what would be required cannot be closed without external methods (axioms, encodings, or borrowed arithmetic).
@@ 10 Formalization structure
All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development (Lean 4,  7,000 LOC; Apache-2.0 licensed) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
@@ 11 Conclusion
This paper presents the KO7 operator-only kernel, proving strong normalization and confluence for the guarded safe subrelation. These results yield a certified normalizer and unique normal forms via Newman’s Lemma. An explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$) is established, blocking confluence for the full relation and motivating the SafeStep fragment. The decidability result explains the clash between single-level Gödel encodings and global termination. The impossibility of additive bumps under duplication is formally verified. Future directions include completing a stratified meta-layer, developing arithmetic encodings within KO7, and mapping the boundary between internally definable measures and external ordinal strength.

### v4

@@ 7 Impossibility results: The Fourteen Failed Methods
    Single-Bit Flags: A boolean flag alone is non-monotone; specific steps (e.g., ‘merge void’) can toggle the flag from 0 to 1.
@@
    Structural Depth: Simple depth counters ($\kappa$) fail on collapsing rules like ‘merge_cancel’ where the result may have equal depth to the input.
@@
    Local Helper Lemmas (Methodological): Attempts to patch specific rules with local lemmas failed due to the global interconnectedness of the recursion.
@@
    Unchecked Recursion: The raw ‘rec_succ’ rule itself serves as a counterexample pattern; without the SafeStep guards, it permits unbounded recursion that is not handled by the audited internal families.
  11. 11.
    Path Orders (Meta-Theoretical): Rely on the universal Subterm Property ($f(t)>t$), which injects an external structural axiom and artificial fixed point not present in the pure calculus.
  12. 12.
    Pure Precedence Orders: Measures relying strictly on head constructor precedence fail on collapsing rules like merge_cancel because the RHS can have the exact same head constructor as the LHS.
  13. 13.
    Linear Weight Functions (KBO-style): Purely additive weight functions fail on the duplication in rec_succ regardless of the constants chosen, as the duplicated argument can grow arbitrarily large.
  14. 14.
    Standard Tree Depth: Measuring the maximum depth of the Abstract Syntax Tree (AST) fails because the duplication of $s$ in rec_succ can strictly increase the overall depth of the term.
Taken together, these fourteen failures provide evidence that the termination barrier is structural: within the audited internal families, the necessary invariants are not captured without the triple-lexicographic measure used for SafeStep.
The Boundary Between Code and Meta-Theory
We distinguish between what is machine-checked in the code and what is argued in the meta-theory regarding Path Orders (Item 11). The Lean artifact (Conjecture_Boundary.lean) formally demonstrates that the isolated components of path orders (pure precedence, Item 12, and linear weights, Item 13) mathematically fail to orient the calculus. In classical proofs, successful path-order arguments rely on the universal Subterm Property ($f(t)>t$). While full LPO is a classical external method that could theoretically orient such systems, doing so imports this property. As argued in Section 8.2, this is a meta-theoretical intervention (via Kruskal-style assumptions) rather than an internally definable proof under Definition 9.1. The code demonstrates pressure toward such an axiom; the text analyzes its scope.
@@
More generally for the duplication pattern studied here: polynomial proofs typically require hardcoded constants (e.g., $M(\mathrm{void})=2$) to satisfy inequalities like $2M(s)>M(s)+1$. Such proofs import external arithmetic properties and impose arbitrary values on operators, violating the principle that operators’ semantics should be defined solely by rewrite rules. Changing the hardcoded constant (e.g., $M(\mathrm{void})=1$) collapses the orientation argument. This is why polynomial interpretations in this setting are treated as outside the intended internal method class.
@@ 8 Decidability of Reachability
Theorem 8.1 (Fixed-target reachability).
In a strongly normalizing and confluent calculus, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
@@
The model assumes a finite first-order rewrite calculus over finite terms. The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with confluence (hence unique normal forms), this reduces to one normalization and an equality test.
@@
Small-term reachability in terminating calculi ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1]. This situates our “normalize and compare” decision procedure for KO7 within the known complexity results.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [4], while termination alone does not guarantee decidability: even flat non-linear calculi have undecidable reachability/joinability/confluence [10].
9 Conjecture: Full Termination Limits in Pure Recursive Calculi (PRC)
Sequential Computation
A Pure Recursive Calculus is one capable of sequential computation: any calculus that can express “do $X$, then do YY” without external conditions.
9.1 The Sequential Nature of KO7 and Recursors
The classification of KO7 as a PRC is not arbitrary; it describes the calculus’s capability to internalize sequential control flow. In flat rewrite calculi (e.g., group axioms), rules rewrite terms to equivalent terms. In a sequential calculus like KO7, the rewrite rules define a computational process.
@@ 9.1 The Sequential Nature of KO7 and Recursors
is not merely a reduction; it is a program. It applies the “step” term $s$ repeatedly, $n$ times. This is functionally identical to the recursor in Gödel’s System T or the step function of primitive recursive arithmetic. Both mechanisms allow a calculus to execute a sequence of operations of unbounded length.
@@
Within this PRC framing, sequential computation requires a mechanism to carry the “step” logic ($s$) into the next iteration. In KO7, this is achieved by duplicating $s$: it appears once in the current application ($\mathrm{app}\,s\dots$) and once in the recursive call ($\mathrm{rec}\Delta\dots s\dots$).
This duplication is the structural barrier. As demonstrated in the Impossibility Results (Section 7) and formally witnessed in the artifact by the theorem rec_succ_additive_barrier, this redistribution of $s$ defeats additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). A calculus that internalizes control flow in this way faces a sharp difficulty when proving full termination with simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
@@
Standard calculi escape this trap by stepping outside the calculus.
  • Proof Assistants (Coq, Lean): Use a meta-logical layer (guard predicates, structural recursion checks) to enforce termination externally. The object calculus does not prove termination by itself; the meta-system checks the required obligations.
@@
KO7’s SafeStep fragment follows this pattern by imposing external guards ($\delta$-flag and $\kappa^{M}$ side conditions). The conjecture below is scoped to Pure Recursive Calculi (Definition 9.2), not to all operator-only kernels.
@@
In this PRC setting, ordered computation is represented with a recursor. The recursor redistributes its step argument across recursive calls. This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
Definition 9.1 (Internally definable measure).
@@
An internally definable termination method for signature $\Sigma$ uses only: DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms. No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
Importantly, an internally definable method must be derived strictly from the operational semantics (the rewrite rules). It cannot inject universal structural axioms from the meta-theory. For example, standard simplification orders (LPO/RPO/MPO) are explicitly excluded because they rely on the universal Subterm Property ($f(t)>t$). As demonstrated by Ferreira and Zantema [8], this property acts as an imported precondition axiom, artificially establishing a fixed point (an anchor) that the pure calculus does not possess. Furthermore, as established in the proof-theoretic literature (e.g., Buchholz [3], Dershowitz & Okada [6]), the well-foundedness of these orders relies on Kruskal’s Tree Theorem, which Gallier [9] notes is a classic independence phenomenon requiring transfinite induction at or above $\varepsilon_{0}$. The distinction is: an internally definable method must be self-contained within the operator language; any proof that borrows structural axioms or well-ordering strength from an external theory is performing meta-level reasoning, not internal reasoning.
Definition 9.2 (Pure Recursive Calculus (PRC)).
A finite first-order rewrite calculus is a Pure Recursive Calculus (PRC) if it satisfies all four of the following conditions:
  1. 1.
    No variable binding. The signature contains no binders; there is no substitution meta-theory and no free variables. All term structure is encoded via constructors, not via variable names.
  2. 2.
    No external memory. The calculus has no external tape, registers, or environment. All computational state is carried within the term itself.
  3. 3.
    No imported axioms. All primitive objects are defined purely as syntactic constructors (e.g., void) governed by the reduction rules, rather than being imported as external mathematical constants with pre-existing axiomatic properties.
  4. 4.
    Unrestricted step-duplicating recursor. The signature contains a recursor rule of the form
    $\mathrm{rec}(b,\,s,\,\sigma(n))\;\to\;f(s,\,\mathrm{rec}(b,\,s,\,n))$
    in which the step argument $s$ appears on both sides and is unrestricted: no guard condition, type annotation, or structural bound is imposed on $s$.
KO7 with full Step is a PRC. Classical Turing machines fail condition 2 (they require an external tape). The untyped lambda calculus fails condition 1 ($\lambda$-abstraction introduces variable binding). Systems of primitive recursive functions fail condition 3 (zero and successor are imported as external primitives, not operator-defined). In the sense of this definition, the class targets a more primitive setting than these classical computational models: the termination barrier identified here arises before arithmetic, variable binding, or external memory enter the picture.
9.2 The Subterm Axiom and Artificial Fixed Points
When classical rewriting theory encounters the duplication barrier, it typically deploys simplification orders (such as LPO or MPO). These orders succeed by enforcing the Subterm Property: the universal rule that any term is strictly greater than its subterms ($f(t)>t$).
However, in the context of a Pure Recursive Calculus, this is a meta-theoretical intervention. While the subterm property is a natural consequence of the inductive definition of terms in the meta-theory, importing that inductive structure into the termination measure of a pure calculus constitutes an external intervention. The pure calculus only “knows” what its reduction rules explicitly define. It does not inherently know that $\mathrm{delta}(t)$ is greater than $t$ unless a rule reduces $\mathrm{delta}(t)$ to $t$. By forcing $f(t)>t$ universally across all constructors, simplification orders impose an external structural constraint on the system. They establish an artificial fixed point (an external anchor or “zero”) that the relational logic never defined.
This suggests a structural limitation: proving termination of a pure sequential calculus may require introducing structural axioms at the meta-level. In this analysis, the anchor is introduced from outside the object calculus.
This structural observation aligns with established proof theory. As demonstrated by Ferreira and Zantema [8], the well-foundedness of path orderings depends on assuming the subterm property as a precondition. Furthermore, in the proof-theoretic literature (e.g., Buchholz [3], Dershowitz & Okada [6]), these orders are formalized as inductively generated relations (least fixed points) whose well-foundedness relies on Kruskal’s Tree Theorem. From the perspective of reverse mathematics, Gallier [9] notes that Kruskal’s theorem is a classic independence phenomenon. As established by Rathjen and Weiermann [12], it requires proof-theoretic strength equivalent to $\Pi^{1}_{1}$-comprehension ($\Pi^{1}_{1}$-CA0) and ordinal bounds reaching into the Veblen hierarchy. Thus, in this framework, orienting a PRC with LPO is treated as meta-level reasoning rather than an internally definable termination proof.
9.3 The Pre-Arithmetic Incompleteness Hierarchy
The conjecture situates KO7 within a three-layer hierarchy of incompleteness results, presented as requiring progressively more external structure:
Table 3: Three layers of computational incompleteness. PRC (Layer 0) requires only an object-level recursor and no external scaffolding; Turing machines (Layer 1) add an external tape and meta-level iteration; Gödel systems (Layer 2) additionally require arithmetic encoding.
Layer Calculus class Minimum requirement Barrier
0 (this work) PRC Order + object-level recursor No known internally definable proof of full termination
1 (Turing) General computation External tape + meta-level transition Cannot decide halting (undecidable)
2 (Gödel) Arithmetic calculus Robinson Q (++ and $\times$) Cannot prove own consistency internally
Why Layer 0 Is Placed Below Layer 1
A classical Turing machine performs sequential computation via a meta-level transition function $\delta:Q\times\Gamma\to Q\times\Gamma\times\{L,R\}$. The iteration (“run for $k$ steps”) is defined outside the machine as a sequence of configuration yields. The tape is external memory; there is no operator within the term language that internalizes the step. In a PRC, by contrast, the recursor $\mathrm{rec}\Delta$ is an object-level constructor: “do the next step” is a rewrite rule applied from inside the term, not a meta-operation on a configuration. This is why a Turing machine fails condition 2 of Definition 9.2 and falls outside the PRC class.
Why Layer 0 Is Placed Below Layer 2
Gödel’s incompleteness theorems require arithmetic sufficient for encoding, at minimum Robinson arithmetic $Q$, with addition, multiplication, zero, and successor as imported primitives. Presburger arithmetic (addition only) and Skolem arithmetic (multiplication only) are both decidable and complete; they escape Gödel because they lack the encoding power of full $Q$. A PRC requires none of this: no imported arithmetic, no Gödel numbering, no self-referential encoding. The termination barrier at Layer 0 arises from structural duplication in the recursor (a constructive phenomenon) rather than from a diagonal self-reference argument. Arithmetic itself is a higher-layer phenomenon: it emerges when the PRC operators are later interpreted (e.g., $\delta$ as successor, $\mathrm{rec}$ as primitive recursion). The Layer 0 barrier exists before that interpretation is imposed. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
Conjecture 9.3 (Scoped Full-Termination Conjecture for Pure Recursive Calculi).
@@ 9.3 The Pre-Arithmetic Incompleteness Hierarchy
For Pure Recursive Calculi (Definition 9.2), no internally definable measure currently known to us proves full-calculus termination. Specifically: let $R$ be an operator-only calculus with a recursor rule of the form
@@
No internally definable measure (Definition 9.1) is currently known to prove termination of such an $R$ when the step argument $s$ is unrestricted. Specifically:
  • Additive and polynomial methods are formally excluded in the audited family. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$; polynomial interpretations without imported arithmetic degrade to additive measures and fail by the same identity (Lemma 4.2). Machine-checked barriers in Meta/Conjecture_Boundary.lean formally witness failure of additive counters, simple lexicographic pairs, flag-only methods, polynomial multiplicative ties, structural depth, and constellation-style hybrids (Section 7).
  • The Subterm Axiom Boundary (LPO/RPO/MPO): Standard path orders are explicitly excluded from internally definable methods because they rely on the universal Subterm Property ($f(t)>t$). This property acts as an imported structural axiom, establishing an anchor that the pure calculus does not define via its own rewrite rules. Forcing $f(t)>t$ universally bypasses the duplication obstruction by adding meta-level structure. Under Definition 9.1, importing this axiom violates the “no imported axioms” boundary for PRCs; therefore, such arguments are treated as meta-theoretical rather than internal proofs.
    Note on orientation layers: The canonical SafeStep termination proof in this paper is the computable stack $(\delta,\kappa^{M},\tau)$ from Meta/ComputableMeasure.lean: rec_zero is discharged by DM on $\kappa^{M}$, rec_succ by the $\delta$ drop, and merge_cancel/eqW tie branches by $\tau$. MPO/LPO appear only as external meta-theoretic comparators in the boundary discussion, not as the internal certified certificate.
Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside this scope.
@@
The conjecture applies to Pure Recursive Calculi (Definition 9.2) under the internal-method boundary above. Calculi that fail any of the four PRC conditions (by introducing external memory, variable binding, imported axioms, or a guarded recursor) may lie outside this scope.
@@
In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, additive measures fail to decrease strictly. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$. The guarded SafeStep fragment terminates because the $\delta\text{-flag}$ restricts allowable step arguments; without such restriction, the full calculus admits unbounded recursive depth. At present, no internally definable ranking-function closure is known in this setting without external methods (axioms, encodings, or borrowed arithmetic).
@@ 10 Formalization structure
All theorem statements currently asserted in the build are formally proven without unproven postulates. Confluence theorems in Meta/Newman_Safe.lean are stated with a global local-join parameter ($\mathsf{locAll}$). The artifact discharges this parameter via locAll_safe from Meta/Confluence_Safe.lean to obtain unconditional confluentSafe and unique normal forms for SafeStep. The formal development (Lean 4, $\sim$6,000 LOC; dual-licensed for academic and commercial use) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
@@ 11 Conclusion
This paper presents the KO7 operator-only kernel, proving strong normalization for a guarded safe subrelation and formalizing a Newman confluence engine. These results yield a certified normalizer; unique normal forms follow unconditionally for the safe fragment. An explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is established for the full relation, motivating the SafeStep fragment. The decidability result explains the clash between single-level Gödel encodings and global termination. The impossibility of additive bumps under duplication is formally verified.